# PaCE: Parsimonious Concept Engineering for Large Language Models

### v1

@@ B.4 Implementation Details
In our experiments, each response of the target LLM is set at a maximum of $512$ tokens. We set the scalar of the representation reading for concept vectors to $3.0$. The experiments are conducted on a workstation of 8 NVIDIA A40 GPUs. Activation vectors are extracted from the last $19$ layers of the target LLM’s decoder layer. For each input prompt, the decomposition is conducted on the inference process of the first next token, and the linear weights are reused for all next token predictions. All LLaMA-2 models in our experiments are the chat version (i.e., optimized for dialogue use cases). GPT-4-0125 is used for dictionary construction and concept partition. All alignment experiments use the top $10,000$ concepts from our PaCE-1M dataset to construct the concept dictionary (as Table 3 validates that the performance is high and does not change much after dictionary size $\geq 9000$). Each concept of PaCE-1M has at least 30 contextual sentences. For each alignment task, PaCE removes the top 50 undesirable concepts ranked by the GPT partitioner (§3). When solving the optimization problem for decomposition in §3.3, we set $\tau=0.95$ and $\alpha=0.05$ following the observations in [82]. The MMLU evaluation is the 5-shot setting where 5 demonstrations are provided during question prompting. After retrieving the relevant knowledge (with the contriever [27]) from Wikipedia for concept synthesis, we take the top-5 ranked facts to append the instruction of LLM. The FAISS-indexed [31] Wikipedia is a snapshot of the 21 million disjoint text blocks from Wikipedia until December 2018.
For the prompting baseline in Table 1 and Table 4, the instruction to the target LLM is to let the model be aware of the partitioned undesirable concepts and not to respond contents relevant to these concept:
@@ B.5 Ablation Study
In this section, we describe the details of the ablation study. In Table 3, we begin with decomposing the input on the five open-sourced [^5] https://github.com/andyzoujm/representation-engineering/tree/main/data/emotions emotion concepts (anger, disgust, fear, happiness, sadness, surprise) [88] and removing only the concept ‘disgust’ with no partitioner (automatic selection of relevant concepts) or clustering (manual selection of relevant concept clusters). Then the design of Decomposition on $10^{4}$ Concepts means that the dictionary is updated to be the top $10,000$ concepts in our PaCE-1M dataset and the concept ‘harmful’ from our dataset is removed. The Clustering of Concepts indicates that we run subspace clustering (detailed in Appendix C.2) and manually choose to remove all concepts of the cluster 125 with the PaCE-solved coefficients: ‘murder’, ‘evil’, ‘kill’, ‘violence’, ‘dirty’, ‘bomb’, ‘violent’, ‘armed’, ‘gross’, ‘savage’, ‘vicious’, ‘explosive’, ‘abuse’, ‘assault’, ‘penetration’, ‘cruelty’, ‘corruption’, ‘tyranny’, ‘tortured’, ‘notorious’, ‘militant’, ‘bloody’, ‘insult’, ‘lure’, ‘ruthless’, ‘inhuman’, and ’brutal’. Concept Partitioner means that we instruct GPT-4 to classify every concept as benign or undesirable (with a ranking score) and remove the top $10$ undesirable concepts with the PaCE-solved weights. Lastly, the Removal of Top $50$ Concepts suggests that we remove the top 50 concepts in the undesirable partition.
Table 3 shows the effect of the dictionary size on three metrics (safety score, response fluency, and the average time per response). The fluency metric remains relatively consistent across different dictionary sizes, showing that PaCE’s decomposition maintains the general linguistic performance. Safety score and response time increase as the dictionary size increases. We observe that the safety performance does not increase too much after the dictionary size changes from 9000 to 10000. This validates our experiment choice of the dictionary size in this interval.
B.6 Limitations, Societal Impacts, and Future Works
@@
Parsimonious Concept Representation. In this paper, we follow the current practice (§2.2) to represent a concept by a single vector. Nonetheless, there might be several alternatives. Results on linear polysemy [2, 15, 84] suggest that a concept might be better represented by multiple vectors or low-dimensional linear subspaces, each corresponding to different semantic meanings. A concept vector may also be sparse, i.e., having a few non-zero entries: the work of [8, 20] identifies some expert neurons in LLMs associated with each concept, and the authors of [40] observe that some layer in a transformer block manifests very sparse activation across all depth levels of various transformer architectures for different tasks. Inspired by how parsimonious structures can be used to accelerate the inference of LLMs [12], controlling the LLMs could also be made faster.
@@ C.2 Subspace Clustering on Concept Vectors
Algorithm. We apply Elastic Net Subspace Clustering (EnSC) on the preprocessed vectors to obtain $200$ clusters. The parameters of EnSC is set to $\tau=1$ and $\gamma=100$.
@@ Appendix D Textual Results
D.1 Baseline Responses

### v2

@@ B.4 Implementation Details
The experiments are conducted on a workstation of 8 NVIDIA A40 GPUs. Each response of the target LLM is set at a maximum of $512$ tokens. Activation vectors are extracted from the last-$29^{\text{th}}$ to the last-$11^{\text{th}}$ layer (totaling $19$ layers) of the target LLM’s decoder layers. All LLaMA-2 models in our experiments are the chat version (i.e., optimized for dialogue use cases).
Concept Dictionary Construction and Partition. We set the scalar of the representation reading for concept vectors to $3.0$. GPT-4-0125 is used for dictionary construction and concept partition. Each concept of PaCE-1M has at least 30 contextual sentences. For each alignment task, PaCE removes the top 50 undesirable concepts ranked by the GPT partitioner (§3). After retrieving the relevant knowledge (with the contriever [27]) from Wikipedia for concept synthesis, we take the top-5 ranked facts to append the instruction of LLM. The FAISS-indexed [31] Wikipedia is a snapshot of the 21 million disjoint text blocks from Wikipedia until December 2018.
Activation Decomposition and Intervention. For each input prompt, the decomposition is conducted on the inference process of the first next token, and the linear weights are reused for all next token predictions. All alignment experiments use the top $10,000$ concepts from our PaCE-1M dataset to construct the concept dictionary (as Table 3 validates that the performance is high and does not change much after dictionary size $\geq 9000$). When solving the optimization problem for decomposition in §3.3, we set $\tau=0.95$ and $\alpha=0.05$ following the observations in [82].
Evaluation. The MMLU evaluation is the 5-shot setting where 5 demonstrations are provided during question prompting. For the prompting baseline in Table 1 and Table 4, the instruction to the target LLM is to let the model be aware of the partitioned undesirable concepts and not to respond contents relevant to these concept:
@@ B.5 Ablation Study
In this section, we describe the details of the ablation study. In Table 3, we begin with decomposing the input on the five open-sourced [^6] https://github.com/andyzoujm/representation-engineering/tree/main/data/emotions emotion concepts (anger, disgust, fear, happiness, sadness, surprise) [88] and removing only the concept ‘disgust’ with no partitioner (automatic selection of relevant concepts) or clustering (manual selection of relevant concept clusters). Then the design of Decomposition on $10^{4}$ Concepts means that the dictionary is updated to be the top $10,000$ concepts in our PaCE-1M dataset and the concept ‘harmful’ from our dataset is removed. The Clustering of Concepts indicates that we run subspace clustering (detailed in Appendix C.2) and manually choose to remove all concepts of the cluster 125 with the PaCE-solved coefficients: ‘murder’, ‘evil’, ‘kill’, ‘violence’, ‘dirty’, ‘bomb’, ‘violent’, ‘armed’, ‘gross’, ‘savage’, ‘vicious’, ‘explosive’, ‘abuse’, ‘assault’, ‘penetration’, ‘cruelty’, ‘corruption’, ‘tyranny’, ‘tortured’, ‘notorious’, ‘militant’, ‘bloody’, ‘insult’, ‘lure’, ‘ruthless’, ‘inhuman’, and ‘brutal’. Concept Partitioner means that we instruct GPT-4 to classify every concept as benign or undesirable (with a ranking score) and remove the top $10$ undesirable concepts with the PaCE-solved weights. Lastly, the Removal of Top $50$ Concepts suggests that we remove the top 50 concepts in the undesirable partition.
Figure 3 shows the effect of the dictionary size on three metrics (safety score, response fluency, and the average time per response). The fluency metric remains relatively consistent across different dictionary sizes, showing that PaCE’s decomposition maintains the general linguistic performance. Safety score and response time increase as the dictionary size increases. We observe that the safety performance does not increase too much after the dictionary size changes from 9000 to 10000. This validates our experiment choice of the dictionary size in this interval.
Figure 11: Ablation study for the regularization $\tau$.
$\tau$ Note Safety (%, $\uparrow$)
0 Pure $\ell_{2}$ 68.9
0.35 N.A. 65.4
0.65 N.A. 71.6
0.95 N.A. 72.0
1.0 Pure $\ell_{1}$ 66.5
Figure 11 shows that the regularization with $\tau=0.95$ yields the best safety performance among the five choices. Pure ridge regression ($\tau=0$) and pure lasso regression ($\tau=1$) do not perform as well as the mixed regularization strategy.
B.6 Providing Dictionary and Scores to Target LLM
For clarity, we elaborate on more details of how PaCE uses the concept dictionary. In our paper, dictionaries are a collection of concept vectors and are frozen for representation decomposition. First, the LLM takes an input prompt (e.g., malicious requests). Then an activation engineering framework [88, 72] extracts the activations at each decoder block of the transformer. Such extraction results in a vector corresponding to the input prompt, which can then be modified for steering in different ways. For PaCE, the steering has two main stages. Stage 1 pre-computes the large-scale concept dictionary offline and the partition (i.e., scores) of which concepts represent benign/harmful concepts. Stage 2 extracts the representation of an input prompt at inference time and uses sparse coding to decompose this as the linear combination of atoms in our frozen dictionary. We then modify this linear combination by removing undesirable components and proceeding with inference in the LLM with the detoxified representation.
B.7 Limitations, Societal Impacts, and Future Works
@@
Parsimonious Concept Representation. In this paper, we follow the current practice (§2.2) to represent a concept by a single vector. Nonetheless, several alternatives could be explored. Results on linear polysemy [2, 15, 84] suggest that a concept might be better represented by multiple vectors or low-dimensional linear subspaces, each corresponding to different semantic meanings. A concept vector may also be sparse, i.e., having a few non-zero entries: the work of [8, 20] identifies some expert neurons in LLMs associated with each concept, and the authors of [39] observe that some layer in a transformer block manifests very sparse activation across all depth levels of various transformer architectures for different tasks. Inspired by how parsimonious structures can be used to accelerate the inference of LLMs [12], controlling the LLMs could also be made faster.
@@ C.2 Subspace Clustering on Concept Vectors
Algorithm. We apply Elastic Net Subspace Clustering (EnSC) [82] on the preprocessed vectors to obtain $200$ clusters. The parameters of EnSC is set to $\tau=1$ and $\gamma=100$.
@@ Appendix D Textual Results
D.1 Baseline Responses and Additional Benchmark
@@
Table 5: Detoxification evaluation for PaCE, representation manipulation, and training-free baselines on AdvBench.
Vanilla PE VecAdd OrthoProj PaCE (Ours)
LlaMA2-7B-Chat 11.72 91.90 94.51 92.81 96.65
LlaMA2-13B-Chat 18.04 93.86 95.33 96.72 99.17
AdvBench [89] adversarially optimizes a jailbreak suffix for a harmful behavior request. Table 5 shows the LlaMA-7B-Chat and LlaMA-13B-Chat safety scores (%, $\uparrow$) on the effective set of suffix attacks for AdvBench harmful behavior set. The detoxification setup follows §4.1. We observe that PaCE outperforms other baselines. We also note that the outperformance of PaCE in §4.1’s jailbreaks is more significant than that in suffix attacks. This is potentially because story-telling and roleplay jailbreaks contain more complex and entangled concepts. Under this scenario, PaCE decomposes the target representation and well estimates the malicious component, while VecAdd and OrthoProj do not model the space sufficiently. In the AdvBench case, instead, the optimized adversarial suffix can be regarded as the text-space inversion of straightforward malicious concepts. PaCE and other defense mechanisms in latent space and prompt space shall effectively defend these suffixes more easily.